OpenStudy (butterflyhope):

Check: Simplify 3a^2 + 4a - a^2 + a; then evaluate for a = 5. My answer: 4a^2 + 4a; 25 Thanks :)

OpenStudy (mathmath333):

its faulty

OpenStudy (mathmath333):

\(\Large\tt \color{black}{3a^2 + 4a - a^2 + a=2a^2+5a=a(2a+5)}\)

OpenStudy (butterflyhope):

Oh, thank you!